determine k so that the line 5x + ky = 32 will pass through the point P(4, 3)
Answer by jim_thompson5910(35256)

5x + ky = 32


5(4) + k(3) = 32 ... Plug in (x,y) = (4,3). Now solve for k.


20 + 3k = 32


3k = 32 - 20


3k = 12


k = 12/3


k = 4